given f(x)= (7/x)-3 find f^-1(x)
Answer by nerdybill(7384)

given f(x)= (7/x)-3 find f^-1(x)
.
f(x)= (7/x)-3
y= (7/x)-3
switch x for y and y for x:
x= (7/y)-3
solve for y:
xy = 7-3y
xy+3y = 7
y(x+3) = 7
y = 7/(x+3)
.
so:
f^-1(x) = 7/(x+3)